# Relocate: A Simple Training-Free Baseline for Visual Query Localization Using Region-Based Representations

### v1

@@ 1 Introduction
Visual Query Localization (VQL) requires localizing the last appearance of an object of interest in a long video. The object of interest is specified via a reference image, also known as the visual query. Figure 1 provides illustrative examples. VQL is an important task, e.g., for surveillance, legal investigations, wildlife monitoring, or simply for tracking down a misplaced item. However, the task presents several unique challenges that push the boundaries of contemporary computer vision methods. For instance, unlike classic object detection models that are trained to identify a fixed set of object categories [32, 5], VQL requires localizing an open-ended range of objects. Additionally, while typical object tracking methods are initialized with a bounding box close to the object’s temporal location in the video [41, 19, 22], the VQL reference image (visual query) often originates outside the video, i.e., there may be no exact or neighboring frame for reliable matching. Further, the object’s appearance may vary significantly from the visual query due to changes in orientation, scale, context, lighting, motion blur, and occlusions. Compounding these issues, the object of interest usually appears briefly in a long, untrimmed video.
Classic VQL methods typically use a stage-wise pipeline [39, 40, 12]: (1) identify all objects in each video frame, (2) compare these objects with the given query to select the most similar ones, and (3) perform bidirectional tracking to obtain a spatio-temporal response. While this pipeline is well-grounded in the object localization literature, it is effective only for large, consistently visible objects that closely match the visual query in short video clips. A more recent work [15] has proposed an end-to-end framework that aims to understand the holistic relationship between a given query and the video, performing spatio-temporal localization in a single step. However, this approach requires extensive training on large, annotated datasets, which can be challenging to obtain. Additionally, since such a method learns a holistic video-query relationship, it must process the entire video for each query, even when multiple queries are to be localized in the same video.
@@
We evaluate Relocate using the Ego4D Visual Query 2D (VQ2D) Localization benchmark [12] and observe significant improvements over prior methods. We achieve a 49% increase in spatio-temporal average precision and a 33% boost in temporal average precision compared to prior state-of-the-art, all without task-specific training.
@@ 2 Related Work
Object Instance Recognition. Early approaches to object instance recognition retrieve images with similar keypoints to the visual query [37, 24, 18, 25]. This works well for objects with distinctive textures, but fails for low-texture, blurry, and highly occluded objects, which are common in the VQL task. Subsequent research shifts towards using features extracted from convolutional neural networks (CNNs) as descriptors for image retrieval [3, 7, 30, 31, 1, 2]. However, these early CNN-based methods suffer when faced with large changes in scale, rotation, or viewpoint. The advent of vision foundation models such as CLIP [28], DINO [6], and DINOv2 [26] has enabled new strategies for instance recognition and retrieval. Several works [4, 35, 9, 34] have leveraged the cross-modal capabilities of CLIP for specialized retrieval tasks. However, these approaches are primarily designed for scenarios where query objects occupy a substantial portion of the target frame, which is not the case in VQL. Recent work on region-based representations [36] shows that pooling DINOv2 features over SAM regions provides effective example-based object category retrieval. In this work, we apply a similar approach for our initial search but propose enhancements to refine, track, and expand the initial query.
@@
Video Object Tracking. Conventional tracking approaches are initialized with a bounding box in an initial frame of the target video, and they track objects through gradual appearance changes. In contrast, VQL often involves queries from frames outside the target video, leading to significant variations in appearance, viewpoint, and context. Nevertheless, tracking remains crucial in VQL. Early tracking approaches relied on motion and appearance cues, using correlation filters and keypoint matching [33, 13, 16]. More recently, trackers leverage large Transformer models to effectively track one or more objects in long videos [8, 11, 21, 19, 23, 41]. Some recent works have also extended SAM [17] for tracking [10, 42, 29]. In this work, we use SAM 2 [29].
@@
Visual Query Localization. The Ego4D benchmark [12] recently introduced VQ2D, a benchmark for VQL. The initial approach, proposed as a baseline in Ego4D, employs a three-stage detection and tracking framework: performing frame-level detection, identifying the most recent detection peak across time, and applying bi-directional tracking to determine the complete temporal extent of the target object [12]. Subsequent works enhance the framework’s performance through various refinements like incorporating negative frame sampling to reduce false positives [39] and leveraging background objects as contextual cues [40]. In this work, we adopt a similar stage-wise framework, introducing key design decisions and refinements that significantly improve the baseline. A more recent effort trains a network to learn the query-video relationship and perform VQL in a single step. However, it relies on a large amount of annotated data, which is costly to obtain. In contrast, we present a training-free solution.
@@ 3 Relocate
We focus on VQL, the task of localizing the last occurrence of a query object in a video. Formally, given a video $\mathcal{V}$, a visual query $\mathcal{Q}$, and a query time $T$, the objective is to predict a response track =$\mathcal{R}=\{b_{s},b_{s+1},\dots,b_{e}\}$ that localizes and tracks the latest occurrence of the query object prior to time $T$. Here, $\mathcal{Q}$ is specified by a bounding box in a reference frame, $s$ and $e\leq T$ denote the first and the last frames in which the query object is visible, and $b_{i}$ represents a bounding box around the object in frame $i$.
@@ 3.1 Prepare
To address VQL, it is imperative to encode and represent all regions/objects in each frame of a video. However, creating patch-level features for each frame results in an excessive number of tokens for long videos, making the task computationally expensive. To address this issue, we propose using region-based representations [36] to encode each frame. As shown in Table 2, the region-based approach significantly reduces the number of tokens needed to encode a video. Moreover, this produces semantically meaningful object tokens, simplifying entity search for a given query.
@@
1. Segment Objects. We first generate a binary mask for each region/object in all frames of the video via a segmentation model. Formally, for a frame $f_{t}\in\mathbb{R}^{H\times W\times 3}$, we extract a set of binary masks $M_{t}=\{m_{t1},m_{t2},\dots\}$, one for each region. Here, $m_{ti}\in\mathbb{R}^{H\times W}$ is a binary mask with 1 in the area occupied by the object $i$ and 0 otherwise. In this work, we use SAM ViT-H [17] as the segmentation model, which processes each frame at a resolution of $1024\times 1024$.
2. Extract Features. We then extract dense features from every frame of the video via a pretrained vision model. Formally, for every frame $f_{t}\in\mathbb{R}^{H\times W\times 3}$ we compute a high-dimensional feature map $h_{t}\in\mathbb{R}^{h\times w\times d}$. In this work, we experiment with different feature extractors and choose DINO ViT-B/8 [6] with a frame resolution of $384\times 512$.
3. Resize and Pool. Subsequently, we resize the frame features $h_{t}\in\mathbb{R}^{h\times w\times d}$ to $\overline{h}_{t}\in\mathbb{R}^{H\times W\times d}$ to fit the height and width of the masks. We then pool the features within each object mask $m_{ti}$ to get an object token $o_{ti}\in\mathbb{R}^{d}$. Following the insights from [36], we use bilinear interpolation to resize the features and average pooling to aggregate the features within the mask.
@@ 3.2 Search
1. Compute Similarity. We compute pairwise cosine similarity between all object tokens and query tokens. Initially, only one query token is generated from the given visual query, but multiple query tokens can be created for the same object. When multiple query tokens are present, we use the maximum similarity score across all tokens as the object score, ensuring that an object is selected if it matches any view of the query. Formally, for the $i^{\text{th}}$ object token in frame $t$ (i.e., $o_{ti}$) and $m$ query tokens $[q_{j}]_{j=1}^{m}$, the similarity score $s_{ti}$ is computed via
$\vspace{-0.5pt}s_{ti}=\max_{j\in\{1,\ldots,m\}}\frac{o_{ti}\cdot q_{j}}{\|o_{ti}\|\|q_{j}\|}.$ (1)
We also studied test-time training (TTT) [38], but it performed worse than a simple cosine similarity search due to an insufficient number of query tokens for effective training.
@@ 4 Experiments
SiamRCNN [12] 0.15 0.22 43.2 32.9
NFM [39] 0.19 0.26 47.9 37.9
CocoFormer [40] 0.19 0.26 47.7 37.7
VQLoC [40] 0.22 0.31 55.9 47.1
@@
SiamRCNN [12] 0.13 0.21 41.6 34.0
CocoFormer [40] 0.18 0.26 48.1 43.2
VQLoC [40] 0.24 0.32 55.9 45.1
@@
Table 1: Results on Ego4D VQ2D benchmark. The validation results are taken from [40], and the test results are obtained from the challenge leaderboard.
@@ 4.1 Evaluation
Dataset. We evaluate Relocate on the Ego4D VQ2D dataset [12], a large collection of egocentric videos annotated for VQL within episodic memory. On average, the videos in this dataset are 140 seconds long, and the target response tracks last for roughly 3 seconds. The dataset comprises 13,600 training, 4,500 validation, and 4,400 test queries, annotated across 262, 87, and 84 hours of video, respectively. We use the validation set for our development and ablations. To our knowledge, VQ2D is the only publicly available dataset for VQL.
@@
Timing Analysis. On average, preparing a 1000-frame video (i.e., extracting object tokens) takes 1422.5 seconds with our current setup. Subsequent steps are faster: search takes 0.8 seconds, refinement 12.6 seconds, tracking 26.3 seconds, and generating additional queries from a response track is instantaneous. Note that our current implementation is not optimized for speed. For applications requiring faster localization, several simple optimizations can be made with minimal impact on the metrics, such as using HNSW [20] for search, increasing the batch size for feature extraction, using faster SAM variants [27] for region generation, and employing faster trackers. We leave these enhancements for future work.
@@ 4.2 Design Decisions
Region vs. Patch Representations. Most contemporary approaches to object localization and tracking encode video frames using patch-based representations, where each frame is divided into non-overlapping patches and encoded with a vision transformer. Region-based representations derived from pretrained vision models [36] have recently been proposed as an efficient alternative. These representations offer both semantic richness and compact encoding—qualities that align with the requirements of VQL.
Notably, region token count depends on the scene content (detected objects or regions) rather than spatial parameters (image dimensions and patch size). As a result, in most practical cases, a region-based approach yields a much lower token count than the patch-based approach. As shown in Table 2, a region-based encoding of a 200-second VQ2D video produces 24×24$\times$24 × fewer tokens compared to a patch-based encoding at an image resolution of $384\times 512$ with a patch size of 8. Moreover, this difference drastically increases if we increase the resolution; e.g. if the image resolution is increased to $1024\times 1024$, a region-based encoding produces 126×126$\times$126 × fewer tokens than a patch-based encoding. Thus, the decoupling of token count from spatial parameters allows the region-based approach to leverage the benefits of smaller patches and high-resolution images without incurring the computational overhead typically associated with increased token counts.
@@
Feature Extraction. To search and retrieve objects that match a given query, VQL requires capturing fine-grained details from each video frame, and the DINO models are particularly well-suited for this task [36, 14]. We compare DINO [6] and DINOv2 [26] using various ViT backbones, with results shown in Table 3. Our findings indicate that DINO ViT-B/8 outperforms ViT-B/16, likely due to its smaller patch size enabling finer feature extraction. Additionally, DINOv2 ViT-L/14 shows stronger frame-level retrieval, and DINO ViT-B/8 is better at spatial localization of objects. We choose DINO ViT-B/8 as our feature extractor due to its higher spatio-temporal localization success rate, smaller model size for faster inference, and reduced feature dimensionality requiring less memory.
@@ Appendix B Evaluating SAM 2 on VQ2D
Jiang et al. [15] demonstrated significant limitations in VQL capabilities among contemporary tracking systems. Specifically, they showed that STARK [41], a state-of-the-art visual tracker at the time, achieves only a 0.04 $\text{stAP}_{25}$ score on the VQ2D validation set. Since then, tracking systems have advanced considerably. To evaluate the capabilities of current tracking systems, we test SAM 2 [29] on the VQL task.

### v2

@@ 1 Introduction
Visual Query Localization (VQL) requires localizing the last appearance of an object of interest in a long video. The object of interest is specified via a reference image, also known as the visual query. Figure 1 provides illustrative examples. VQL is an important task, e.g., for surveillance, legal investigations, wildlife monitoring, or simply for tracking down a misplaced item. However, the task presents several unique challenges that push the boundaries of contemporary computer vision methods. For instance, unlike classic object detection models that are trained to identify a fixed set of object categories [Ren2015FasterRT, Carion2020EndtoEndOD], VQL requires localizing an open-ended range of objects. Additionally, while typical object tracking methods are initialized with a bounding box close to the object’s temporal location in the video [Yan2021LearningST, Ma2022UnifiedTT, Mayer2022BeyondST], the VQL reference image (visual query) often originates outside the video, i.e., there may be no exact or neighboring frame for reliable matching. Further, the object’s appearance may vary significantly from the visual query due to changes in orientation, scale, context, lighting, motion blur, and occlusions. Compounding these issues, the object of interest usually appears briefly in a long, untrimmed video.
Classic VQL methods typically use a stage-wise pipeline [Xu2022NegativeFM, Xu2022WhereIM, Grauman2021Ego4DAT]: (1) identify all objects in each video frame, (2) compare these objects with the given query to select the most similar ones, and (3) perform bidirectional tracking to obtain a spatio-temporal response. While this pipeline is well-grounded in the object localization literature, it is effective only for large, consistently visible objects that closely match the visual query in short video clips. A more recent work [Jiang2023SingleStageVQ] has proposed an end-to-end framework that aims to understand the holistic relationship between a given query and the video, performing spatio-temporal localization in a single step. However, this approach requires extensive training on large, annotated datasets, which can be challenging to obtain. Additionally, since such a method learns a holistic video-query relationship, it must process the entire video for each query, even when multiple queries are to be localized in the same video.
@@
We evaluate Relocate using the Ego4D Visual Query 2D (VQ2D) Localization benchmark [Grauman2021Ego4DAT] and observe significant improvements over prior methods. We achieve a 49% increase in spatio-temporal average precision and a 33% boost in temporal average precision compared to prior state-of-the-art, all without task-specific training.
@@ 2 Related Work
Object Instance Recognition. Early approaches to object instance recognition retrieve images with similar keypoints to the visual query [Sivic2003VideoGA, Nistr2006ScalableRW, Lowe2004DistinctiveIF, Obdrzlek2005SublinearIF]. This works well for objects with distinctive textures, but fails for low-texture, blurry, and highly occluded objects, which are common in the VQL task. Subsequent research shifts towards using features extracted from convolutional neural networks (CNNs) as descriptors for image retrieval [Babenko2014NeuralCF, Chandrasekhar2016APractical, Razavian2014CNNFO, Razavian2014ABF, Azizpour2014FromGT, Babenko2015AggregatingLD]. However, these early CNN-based methods suffer when faced with large changes in scale, rotation, or viewpoint. The advent of vision foundation models such as CLIP [Radford2021LearningTV], DINO [Caron2021EmergingPI], and DINOv2 [Oquab2023DINOv2LR] has enabled new strategies for instance recognition and retrieval. Several works [Baldrati2022EffectiveCA, Saito2023Pic2WordMP, Chen2020LearningJV, Sain2023CLIPFA] have leveraged the cross-modal capabilities of CLIP for specialized retrieval tasks. However, these approaches are primarily designed for scenarios where query objects occupy a substantial portion of the target frame, which is not the case in VQL. Recent work on region-based representations [ShlapentokhRothman2024RegionBasedRR] shows that pooling DINOv2 features over SAM regions provides effective example-based object category retrieval. In this work, we apply a similar approach for our initial search but propose enhancements to refine, track, and expand the initial query.
@@
Video Object Tracking. Conventional tracking approaches are initialized with a bounding box in an initial frame of the target video, and they track objects through gradual appearance changes. In contrast, VQL often involves queries from frames outside the target video, leading to significant variations in appearance, viewpoint, and context. Nevertheless, tracking remains crucial in VQL. Early tracking approaches relied on motion and appearance cues, using correlation filters and keypoint matching [Ross2008IncrementalLF, Henriques2014HighSpeedTW, Kalal2012TrackingLearningDetection]. More recently, trackers leverage large Transformer models to effectively track one or more objects in long videos [Chen2021TransformerT, Cui2022MixFormerET, Mayer2022TransformingMP, Ma2022UnifiedTT, Meinhardt2021TrackFormerMT, Yan2021LearningST]. Some recent works have also extended SAM [Kirillov2023SegmentA] for tracking [Cheng2023SegmentAT, Yang2023TrackAS, Ravi2024SAM2S]. In this work, we use SAM 2 [Ravi2024SAM2S].
@@
Visual Query Localization. The Ego4D benchmark [Grauman2021Ego4DAT] recently introduced VQ2D, a benchmark for VQL. The initial approach, proposed as a baseline in Ego4D, employs a three-stage detection and tracking framework: performing frame-level detection, identifying the most recent detection peak across time, and applying bi-directional tracking to determine the complete temporal extent of the target object [Grauman2021Ego4DAT]. Subsequent works enhance the framework’s performance through various refinements like incorporating negative frame sampling to reduce false positives [Xu2022NegativeFM] and leveraging background objects as contextual cues [Xu2022WhereIM]. In this work, we adopt a similar stage-wise framework, introducing key design decisions and refinements that significantly improve the baseline. A more recent effort trains a network to learn the query-video relationship and perform VQL in a single step. However, it relies on a large amount of annotated data, which is costly to obtain. In contrast, we present a training-free solution.
@@ 3 Relocate
We focus on VQL, the task of localizing the last occurrence of a query object in a video. Formally, given a video $\mathcal{V}$, a visual query $\mathcal{Q}$, and a query time $T$, the objective is to predict a response track $\mathcal{R}=\{b_{s},b_{s+1},\dots,b_{e}\}$ that localizes and tracks the latest occurrence of the query object prior to time $T$. Here, $\mathcal{Q}$ is specified by a bounding box in a reference frame, $s$ and $e\leq T$ denote the first and the last frames in which the query object is visible, and $b_{i}$ represents a bounding box around the object in frame $i$.
@@ 3.1 Prepare
To address VQL, it is imperative to encode and represent all regions/objects in each frame of a video. However, creating patch-level features for each frame results in an excessive number of tokens for long videos, making the task computationally expensive. To address this issue, we propose using region-based representations [ShlapentokhRothman2024RegionBasedRR] to encode each frame. As shown in Table 2, the region-based approach significantly reduces the number of tokens needed to encode a video. Moreover, this produces semantically meaningful object tokens, simplifying entity search for a given query.
@@
1. Segment Objects. We first generate a binary mask for each region/object in all frames of the video via a segmentation model. Formally, for a frame $f_{t}\in\mathbb{R}^{H\times W\times 3}$, we extract a set of binary masks $M_{t}=\{m_{t1},m_{t2},\dots\}$, one for each region. Here, $m_{ti}\in\mathbb{R}^{H\times W}$ is a binary mask with 1 in the area occupied by the object $i$ and 0 otherwise. In this work, we use SAM ViT-H [Kirillov2023SegmentA] as the segmentation model, which processes each frame at a resolution of $1024\times 1024$.
2. Extract Features. We then extract dense features from every frame of the video via a pretrained vision model. Formally, for every frame $f_{t}\in\mathbb{R}^{H\times W\times 3}$ we compute a high-dimensional feature map $h_{t}\in\mathbb{R}^{h\times w\times d}$. In this work, we experiment with different feature extractors and choose DINO ViT-B/8 [Caron2021EmergingPI] with a frame resolution of $384\times 512$.
3. Resize and Pool. Subsequently, we resize the frame features $h_{t}\in\mathbb{R}^{h\times w\times d}$ to $\overline{h}_{t}\in\mathbb{R}^{H\times W\times d}$ to fit the height and width of the masks. We then pool the features within each object mask $m_{ti}$ to get an object token $o_{ti}\in\mathbb{R}^{d}$. Following the insights from [ShlapentokhRothman2024RegionBasedRR], we use bilinear interpolation to resize the features and average pooling to aggregate the features within the mask.
@@ 3.2 Search
1. Compute Similarity. We compute pairwise cosine similarity between all object tokens and query tokens. Initially, only one query token is generated from the given visual query, but multiple query tokens can be created for the same object. When multiple query tokens are present, we use the maximum similarity score across all tokens as the object score, ensuring that an object is selected if it matches any view of the query. Formally, for the $i^{\text{th}}$ object token in frame $t$ (i.e., $o_{ti}$) and $m$ query tokens $q_{j}]_{j=1}^{m}$, the similarity score $s_{ti}$ is computed via
$\vskip-0.5pts_{ti}=\max_{j\in\{1,\ldots,m\}}\frac{o_{ti}\cdot q_{j}}{\|o_{ti}\|\|q_{j}\|}.$ (1)
We also studied test-time training (TTT) [Sun2019TestTimeTW], but it performed worse than a simple cosine similarity search due to an insufficient number of query tokens for effective training.
@@ 4 Experiments
SiamRCNN [Grauman2021Ego4DAT] 0.15 0.22 43.2 32.9
NFM [Xu2022NegativeFM] 0.19 0.26 47.9 37.9
CocoFormer [Xu2022WhereIM] 0.19 0.26 47.7 37.7
VQLoC [Xu2022WhereIM] 0.22 0.31 55.9 47.1
@@
SiamRCNN [Grauman2021Ego4DAT] 0.13 0.21 41.6 34.0
CocoFormer [Xu2022WhereIM] 0.18 0.26 48.1 43.2
VQLoC [Xu2022WhereIM] 0.24 0.32 55.9 45.1
@@
Table 1: Results on Ego4D VQ2D benchmark. The validation results are taken from [Xu2022WhereIM], and the test results are obtained from the challenge leaderboard.
@@ 4.1 Evaluation
Dataset. We evaluate Relocate on the Ego4D VQ2D dataset [Grauman2021Ego4DAT], a large collection of egocentric videos annotated for VQL within episodic memory. On average, the videos in this dataset are 140 seconds long, and the target response tracks last for roughly 3 seconds. The dataset comprises 13,600 training, 4,500 validation, and 4,400 test queries, annotated across 262, 87, and 84 hours of video, respectively. We use the validation set for our development and ablations. To our knowledge, VQ2D is the only publicly available dataset for VQL.
@@
Timing Analysis. On average, preparing a 1000-frame video (i.e., extracting object tokens) takes 1422.5 seconds with our current setup. Subsequent steps are faster: search takes 0.8 seconds, refinement 12.6 seconds, tracking 26.3 seconds, and generating additional queries from a response track is instantaneous. Note that our current implementation is not optimized for speed. For applications requiring faster localization, several simple optimizations can be made with minimal impact on the metrics, such as using HNSW [Malkov2016EfficientAR] for search, increasing the batch size for feature extraction, using faster SAM variants [Pytorch2023Segment] for region generation, and employing faster trackers. We leave these enhancements for future work.
@@ 4.2 Design Decisions
Region vs. Patch Representations. Most contemporary approaches to object localization and tracking encode video frames using patch-based representations, where each frame is divided into non-overlapping patches and encoded with a vision transformer. Region-based representations derived from pretrained vision models [ShlapentokhRothman2024RegionBasedRR] have recently been proposed as an efficient alternative. These representations offer both semantic richness and compact encoding—qualities that align with the requirements of VQL.
Notably, region token count depends on the scene content (detected objects or regions) rather than spatial parameters (image dimensions and patch size). As a result, in most practical cases, a region-based approach yields a much lower token count than the patch-based approach. As shown in Table 2, a region-based encoding of a 200-second VQ2D video produces $24\times$ fewer tokens compared to a patch-based encoding at an image resolution of $384\times 512$ with a patch size of 8. Moreover, this difference drastically increases if we increase the resolution; e.g. if the image resolution is increased to $1024\times 1024$, a region-based encoding produces $126\times$ fewer tokens than a patch-based encoding. Thus, the decoupling of token count from spatial parameters allows the region-based approach to leverage the benefits of smaller patches and high-resolution images without incurring the computational overhead typically associated with increased token counts.
@@
Feature Extraction. To search and retrieve objects that match a given query, VQL requires capturing fine-grained details from each video frame, and the DINO models are particularly well-suited for this task [ShlapentokhRothman2024RegionBasedRR, Jiang2023FromCT]. We compare DINO [Caron2021EmergingPI] and DINOv2 [Oquab2023DINOv2LR] using various ViT backbones, with results shown in Table 3. Our findings indicate that DINO ViT-B/8 outperforms ViT-B/16, likely due to its smaller patch size enabling finer feature extraction. Additionally, DINOv2 ViT-L/14 shows stronger frame-level retrieval, and DINO ViT-B/8 is better at spatial localization of objects. We choose DINO ViT-B/8 as our feature extractor due to its higher spatio-temporal localization success rate, smaller model size for faster inference, and reduced feature dimensionality requiring less memory.
@@ 5 Conclusion
Acknowledgements
This work is supported in part by ONR N00014-23-1-2383, DARPA HR0011-23-9-0060, NSF 2008387, NSF 2045586, NSF 2106825, NSF MRI 1725729, and NIFA award 2020-67021-32799. We used GPUs at NCSA Delta through allocation CIS240541 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. The views and conclusions expressed are those of the authors, and not necessarily representative of the US Government or its agencies.
[…]
@@ Appendix B Evaluating SAM 2 on VQ2D
Jiang2023SingleStageVQ demonstrated significant limitations in VQL capabilities among contemporary tracking systems. Specifically, they showed that STARK [Yan2021LearningST], a state-of-the-art visual tracker at the time, achieves only a 0.04 $\text{stAP}_{25}$ score on the VQ2D validation set. Since then, tracking systems have advanced considerably. To evaluate the capabilities of current tracking systems, we test SAM 2 [Ravi2024SAM2S] on the VQL task.